OpenStudy (anonymous):

what is a representation of coordinate (75,50)

OpenStudy (anonymous):

Itss about the polar coordinates so somewhat like the equivalent coordinate for 75, 50 degree)

hero (hero):

Yes, you should have mentioned POLAR. That would have been very helpful

OpenStudy (anonymous):

\[(x,y) -> (r,\theta)\] \[r = \sqrt{x^2 + y^2} \] \[\theta = \tan^{-1} (y/x)\] subbing you should get (90.139, 33.690 [deg])

OpenStudy (anonymous):

Sorry for some reason people have been giving me the wrong way of doing the problem